OpenStudy (imtiaz7):

If alpha and beta are the roots of eq. x^2-4x+ 16=0, Find alpha^2 + beta^2 @MelissaHolmes @sarah786

OpenStudy (perl):

the solutions are 2+(2*i)*sqrt(3), 2-(2*i)*sqrt(3)

hartnn (hartnn):

do you know how you can get sum of roots and product of roots of a quadratic equation ?

hartnn (hartnn):

for \(ax^2+bx+c=0\) sum of roots = \(\dfrac{-b}{a}\) and product of roots = \(\dfrac{c}{a}\)

OpenStudy (perl):

alpha^2 + beta^2 = -16

OpenStudy (perl):

remember complex solutions come in conjugate pairs (a+bi) , ( a - bi) and (a+bi)^2 + ( a - bi)^2 = a^2+ 2abi - b^2 + a^2 - 2abi - b^2 = 2a^2 - 2b^2

hartnn (hartnn):

when you have a+b and ab, use this : \(a^2+b^2 = (a+b)^2-2ab\)